### v1

# Horizon-scale effects in astrophysical black holes from mass-dependent regulators

Abstract
Typically, regular black holes are described by a regulator $\ell$ that is much smaller than the horizon radius $r_{\text{h}}$; taking the regulator scale large removes the black hole horizon, leaving a compact, horizonless object. This seemingly justifies the well-known lore that horizon scale effects for astrophysical black holes are suppressed by the dimensionless ratio $\ell/r_{\text{h}}\ll 1$. We show explicitly that regular black hole metrics with $\sim 50\%$ corrections near the horizon are, in fact, possible, without prohibiting a range in black hole masses that would conflict with observation. This opens the doors to potentially observable horizon-scale effects for the light deflection and the shadow around such regular, astrophysical black holes.
Introduction.—Astrophysical black holes are emerging as increasingly relevant testing grounds of gravitational physics [1, 2, 3, 4, 5, 6]. Under some assumptions on the energy and pressure of matter, general relativity predicts that such astrophysical black holes—formed during the late stages of a collapsing massive star—necessarily contain singularities [7, 8, 9]. These singularities show up as divergent gravitational tidal forces in the black hole interior, and prevent a complete description of physics in this regime [10]. This is the black hole singularity problem, a major unresolved problem in gravitational physics. Within general relativity, these singularities are evident in the Schwarzschild metric: an exact, spherically symmetric vacuum solution of the Einstein equations, taking the form [11, 12]
$\begin{split}\mbox{d}s^{2}&=-F(r)\,\mbox{d}t^{2}+\frac{\mbox{d}r^{2}}{F(r)}+r^{2}\mbox{d}\theta^{2}+r^{2}\sin^{2}\theta\,\mbox{d}\varphi^{2}\,,\\ F(r)&=1-\frac{2GM}{r}\,.\end{split}$ (1)
Here, $M$ is the black hole mass, $G$ is the gravitational constant, and we set the speed of light to unity ($c=1$). The event horizon is located at $r_{\text{h}}=2GM$, and the singularity is situated at $r=0$, as can be seen from the Kretschmann invariant $R{}_{\mu\nu\rho\sigma}R{}^{\mu\nu\rho\sigma}=48(GM)^{2}/r^{6}$. It is believed that quantum-gravitational effects change the form of the metric in the black hole interior at $r\ll r_{\text{h}}$, eventually leading to an avoidance of the singularity. In the recent years, a variety of regular black hole metrics have been proposed [13, 14, 15, 16, 17]. Typically, they feature a modification of an otherwise singular metric, parametrized by a length scale $\ell>0$, that removes the singularity at $r=0$. These metrics are not solutions of a fundamental gravitational theory, but can help constrain the leading order observational consequences of regular black holes. A characteristic example is the Hayward metric [16],
$F_{\text{H}}(r)$ $=1-\frac{2GM}{r}\frac{r^{3}}{r^{3}+L^{3}}\,,\quad L^{3}=2GM\ell^{2}\,.$ (2)
Figure 1: A stark difference: The metric function $F(r)$ over the dimensionless distance $r/\ell$ for $\ell=30\,\text{km}$ for two cases of $M$, assuming Eq. (7). Left: $M=30\,M_{\odot}$, the new metric is indistinguishable from the Hayward metric. Right: $M=3M_{\odot}$. The new metric features a horizon and deviates macroscopically from the Schwarzschild metric at the horizon by roughly 8%; the Hayward metric does not describe a black hole, but a solar-mass scale horizonless object, so far unobserved in Nature.
This metric only describes a black hole if it has an outer horizon located at $F_{\text{H}}(r_{\text{h}})=0$, which requires that $\ell<4GM/(3\sqrt{3})$. The regulator $\ell$ is always smaller than the horizon scale, $\ell<r_{\text{h}}$, if a horizon exists. Conversely, a given regulator scale $\ell$ prevents black holes with a mass below $\sim\ell/G$ [^1] This remains true in dynamical black hole formation [18]. This is referred to as a “mass gap.” Tabletop Eötvös-type experiments can constrain the regulator $\ell$, confirming the Newtonian inverse-square law of gravitation to distances of $\sim 50\,\mu\text{m}$ for test masses of $\mathcal{O}(1\,\text{kg})$. A measurement uncertainty of around $0.5\,\mu\text{m}$ [19] bounds $\ell$ in Eq. (2) to
$\ell\lesssim\mathcal{O}(700\,\text{km})\,.$ (3)
Such large values for $\ell$ are possible since for tabletop masses $2GM/c^{2}\sim 10^{-27}\,\text{m}$, and only $L$ has to be $\lesssim 10\mu\text{m}$. The mass-dependent regulator $L$ removes the mass dependence of the function (2) at small distances $r$, which is known as the limiting curvature condition [20, 21, 22], obeyed by most regular black hole metrics [17]. Since $L$ depends on mass, it will be much larger for astrophysical objects. However, the possibility of observing significant horizon-scale effects in astrophysical regular black holes must overcome two immediate challenges:
  • A)
    Theoretical prejudice: it is usually assumed that $\ell$ is $\mathcal{O}(\ell_{\text{Planck}})$, so that effects at the horizon are minuscule and unobservable.
  • B)
    If $\ell$ is large enough so that horizon-scale effects are maximized for a black hole of mass $M$, then the mass gap prohibits black holes with masses less than $M$, in conflict with observation.
In the absence of a model for the origin of Eq. (2), it is not unreasonable to treat $\ell$ as a free parameter and study the phenomenological consequences. However, challenge B cannot be dispensed with so easily and motivates consideration of modifications of the metric. In this Letter, we propose, for the first time, regular black hole metrics that can feature $\mathcal{O}(50\%)$ corrections at the horizon scale in the mass range of astrophysical black holes, without the problems related to the existence of the mass gap. Unlike the regular black hole metrics described in the literature, the class of regular black holes described here allow the existence of black holes in a regime that is most promising for observational tests of modified gravity. For example, the location of the innermost circular orbit for light (the “photon sphere”), gravitational lensing, and the black hole shadow, are all susceptible to horizon-scale deviations in the astrophysical black hole mass range.
Modified metric.—We take the Hayward metric (2) as a starting point, and aim to explore its dependence on the regulator scale $\ell$ as well as its mass parameter $M$ via the dimensionless combination
^$\hat{\ell}=\frac{\ell}{2GM}\,.$ (4)
While ^$\hat{\ell}\ll 1$ is typically assumed in the study of regular black holes, our focus will be on the regime ^$\hat{\ell}\gtrsim 1$ where one has strong deviations from the Schwarzschild form away from the origin. We parametrize
$F(r)$ $=1-\frac{2GM}{r}\frac{r^{3}}{r^{3}+L^{3}}\,,\quad L^{3}=2GM\ell^{2}f(\hat{\ell})\,,$ (5)
where the asymptotic ADM mass of this metric is still given by $M$, but $f(\hat{\ell})>0$ introduces a new regulator $L$ with an unconventional mass dependence. The black hole horizon, in terms of $\hat{r}\equiv r/(2GM)$, lies at $\hat{r}_{\text{h}}^{3}-\hat{r}_{\text{h}}^{2}=-\hat{\ell}^{2}f(\hat{\ell})$. This equation has positive solutions if
^$\hat{\ell}^{2}f(\hat{\ell})\equiv\hat{L}^{3}\leq\frac{4}{27}\,,$ (6)
which one may think of as the “black hole condition.” The mass-dependence of this condition leads to strikingly different results compared to the Hayward metric. Namely, non-monotonic functions $\ell^{2}f(\hat{\ell})$ that increase and decrease at intermediate values for $\hat{\ell}$ generate entirely new, astrophysically viable branches for regular black hole metrics that feature both a horizon and a large value of $\ell$. To illustrate this, the reader may verify that the following function,
$f_{\text{example}}=\frac{1}{1+\hat{\ell}^{4}}\,,$ (7)
provides such an example. The black hole condition $\hat{L}^{3}\leq 4/27$ is guaranteed for $GM\gtrsim 1.284\,\ell$, qualitatively similar to the Hayward case. However, black holes also exist for $GM\lesssim 0.195\,\ell$—an entirely new branch. The exciting property of this branch lies in the fact that it constitutes an upper bound on black hole mass. It allows black holes with masses that are less than the regulator scale (in units where $c=G=1$). This provides a means of addressing challenge B, and opens the door to astrophysically large effects without sacrificing the existence of a horizon. The choice of function $f(\hat{\ell})$ defines a new family of regular black holes, as we discuss in some more detail below.
The impact of this new, mass-dependent regulator can best be captured by estimating horizon-scale effects. Subject to the condition (6), the horizon is located at
$\hat{r}_{\text{h}}=\frac{1}{3}\left\{1+2\cos\left[\tfrac{1}{3}\arccos\left(1-\tfrac{27}{2}\hat{L}^{3}\right)\right]\right\}\,.$ (8)
At the horizon, the deviation of this class of metrics from the Schwarzschild metric is governed by the ratio
$\delta\equiv\left(\frac{\hat{L}}{\hat{r}_{\text{h}}}\right)^{3}~{}\in~{}[0,0.5]\,.$ (9)
The deviation is monotonically increasing and assumes its maximum value of $50\%$ at $\hat{L}^{3}=4/27$. There, usual regular black hole metrics run into challenge B. The function $f(\hat{\ell})$ in the mass-dependent regulator $\hat{L}$ changes the allowed mass ranges for black holes, so that $\mathcal{O}(50\%)$ horizon-scale effects for astrophysical black holes are allowed; Fig. 1 illustrates the effect of the function $f(\hat{\ell})$. Let us now constrain its properties a bit more:
  • For a vanishing regulator at fixed mass $M$ (^$\hat{\ell}\rightarrow 0$) we should recover the Schwarzschild metric. Assuming that $f(\hat{\ell})$ is regular at the origin, this constrains $f\sim\hat{\ell}^{p}$ with $p>-2$ at ^$\hat{\ell}=0$.
  • The limiting curvature condition demands $f$ to be a universal constant. If we merely want to prevent trans-Planckian curvatures, we can relax this condition to only hold for large masses $M$ (at fixed $\ell$), and hence $f\sim\hat{\ell}^{p}$ with $p\geq 0$ for ^$\hat{\ell}\rightarrow 0$.
  • Tabletop experiments for small masses probe the region $M\rightarrow 0$, corresponding to $\hat{\ell}\rightarrow\infty$ at fixed $\ell$. To avoid a more stringent bound than given in Eq. (3), we assume that $f\lesssim 1$ in this limit.
We propose the following, rather general parametrization that captures this essence (see Fig. 2 for a few cases):
$f(\hat{\ell})=\frac{1}{1+a\hat{\ell}^{p}}+\frac{b}{(1+\hat{\ell})\hat{\ell}^{q}}\,.$ (10)
Figure 2: Possible parametrizations of the mass-dependent regulator $\hat{L}^{3}\equiv\hat{\ell}^{2}f(\hat{\ell})$. The black hole condition (6) is satisfied whenever the curves are below the dashed line. Given a regulator $\ell$, this implies a band structure for black hole existence.
In the black hole domain, as described by Eq. (6), the modification is strongest wherever $\hat{L}^{3}$ reaches its maximum value of $4/27$. For example, in the case where $a=1$, $p=4$ and $b=0$, i.e., Eq. (7), maximal effects are obtained for black hole masses near the boundaries of the allowed regions, $\sim 0.195\ell/G$ and $\sim 1.284\ell/G$, while allowing black holes to exist in both intervals. For $\ell\sim 30\,\text{km}$, these boundaries correspond to $\sim 4$ and $\sim 26$ solar masses, respectively. For realistic astrophysical studies, the forbidden mass interval in between can be modified by a different choice of the function $f(\hat{\ell})$. Note that regular supermassive black holes would not be affected by the modification to the metric. Compared to common regular black hole models, the additional stellar black hole “branches” with substantial horizon-scale effects may be ideal targets for direct astrophysical observation.
Figure 3: Mass-dependent deflection angle $\varphi$ (normalized to the Schwarzschild value) for the impact parameter $b=6GM$, as a function of mass $M$, assuming Eq. (7). The bands show the allowed black hole mass ranges.
Consequences.—To demonstrate the qualitatively new features of the metrics described in this Letter, let us set
$\ell\approx 30\,\text{km}$ (11)
and explore observational consequences for astrophysical, solar-mass black holes. For simplicity, we will assume a mass-dependent regulator given by $\{a=1,p=4,b=0\}$.
Light propagation in the metric (5) is described by
$\dot{r}^{2}=E^{2}-V_{\text{eff}}(r)\,,\quad V_{\text{eff}}(r)=\frac{J^{2}}{r^{2}}F(r)\,,$ (12)
where $E\equiv-g_{tt}\dot{t}$ and $J\equiv g_{\varphi\varphi}\dot{\varphi}$ are constants of motion related to the impact parameter $b$ via $b=J/E$, and the dots denote differentiation with respect to the affine parameter of the geodesic $x{}^{\mu}(\lambda)$.
The effective potential $V_{\text{eff}}(r)$ has a maximum outside the black hole horizon at $r=r_{\gamma}$, indicating an unstable circular photon orbit, called the “photon sphere.” The regulator $f(\hat{\ell})$ pushes it inwards, away from $3GM$:
$\hat{L}^{3}=\left(\sqrt{\frac{3GM}{r_{\gamma}}}-1\right)\bigg{(}\frac{r_{\gamma}}{2GM}\bigg{)}^{3}\geq 0\,.$ (13)
The effect is largest for $\hat{L}^{3}=4/27$, resulting in $r_{\gamma}\approx 2.65\,GM$. This is a 12% deviation from the Schwarzschild location $3GM$. A similar statement holds for light deflection around such black holes. Following (13), the deflection angle is
$\varphi(b)=2b\!\!\!\int\limits_{r_{0}(b)}^{\infty}\frac{\mbox{d}r}{r^{2}}\frac{1}{\sqrt{1-\frac{b^{2}}{r^{2}}F(r)}}-\pi\,,$ (14)
where $r_{0}$ denotes the point of closest approach, related to the impact parameter $b$ via $F(r_{0})\,b^{2}=r_{0}^{2}$. We can now compare the light deflection around objects of the same mass $M$, described by different metrics. For definiteness, we fix the impact parameter to be twice the location of the Schwarzschild photon sphere. The maximum resulting effect is of $\mathcal{O}(5\%)$, and is displayed in Fig. 3.
We can now define the “black hole shadow” as the impact parameter for which the bending of light angle grows to infinity—the shadow is the boundary between trapped and deflected light, and hence corresponds to the visible size of a black hole in the sky. It is given by
$b_{\gamma}=\frac{r_{\gamma}}{\sqrt{F(r_{\gamma})}}=\frac{r_{\gamma}}{\sqrt{1-\sqrt{\frac{4GM}{3r_{\gamma}}}}}\,,$ (15)
and the maximum possible deviation from Schwarzschild is again of $\mathcal{O}(5\%)$.
Conclusions.—We have described a new class of regular black hole metrics that can feature potentially observable, astrophysically relevant horizon-scale deviations from the black holes described by general relativity at the percent level. These include, but are probably not limited to, a reduction in the black hole’s apparent size, a smaller photon sphere, and weaker lensing. These effects are made possible across astrophysical black hole mass ranges via a mass-dependent regulator $f(\hat{\ell})$.
Percent-level effects for astrophysical black holes are a novel phenomenon in the field of regular black holes. Usually, large effects are parametrized by a regulator $\ell$ that is thought to be of quantum-gravitational origin, and hence of the order of the Planck length. Effects for astrophysical black holes are thus minuscule, $\ell_{\text{Planck}}/(2GM_{\odot})\sim 10^{-38}$. One may ask: what happens if the fundamental regulator, for some reason, is macroscopic? Then, regular black hole metrics cease to describe black holes with masses below the scale given by the regulator. This is sometimes referred to as a “mass gap,” and leads to unwanted consequences when the regulator is macroscopic.
In this Letter, we resolved this problem by exploring the option of a large regulator $\ell$ that is potentially of the order of kilometers, and constructed a regular black hole metric that still features a horizon for masses at or below $\sim\ell/G$. This is made possible by a mass-dependent regulator
$\hat{L}^{3}=2GM\ell^{2}f\left(\frac{\ell}{2GM}\right)\,,$ (16)
where the function $f$ is a new ingredient. Even in the case of $f=1$ the regulator is mass-dependent [16], mandated by the limiting-curvature condition. It seems like a natural extension to allow for a more general interplay between $L$ and $M$. As we have shown, if this function is non-monotonic in its dimensionless argument, the black hole mass spectrum is enlarged. Previously excluded mass ranges are now accessible and permit black hole geometries featuring horizons, while other ranges remain inaccessible. In a simple case, cf. Eq. (7), we find
$GM\gtrsim 1.284\ell\,,\quad GM\lesssim 0.195\,\ell\,.$ (17)
While the first constraint constitutes a mass gap, the second condition is qualitatively new. This, to the best of our knowledge, is the first example of a “black hole band spectrum” in the context of regular black holes.
One may wonder about the dynamical aspects of these black holes and their non-trivial band structures. For example, is it possible to have two black holes from the low-mass part of the spectrum, make them collide, and end up in the regime where no black holes are allowed? And, in that case, what is the nature of the horizonless end product after a collision of these black holes? It is possible that instabilities of this compact, horizonless object lead to emission of gravitational waves that only terminates once the object is again inside the lower mass range and a horizon has formed. In the presence of more involved regulators $f(\hat{\ell})$ one could have multiple, disconnected bands for these types of black holes, giving rise to a multitude of transitions between bands, accompanied by a characteristic amount of gravitational wave emission. While certainly interesting, we will leave dynamical aspects to future work.
We also briefly comment on the issue of supermassive black holes. Since their mass scale exceeds that of astrophysical black holes up to a power of $10^{9}$ [23], the effects described in this Letter are suppressed by roughly this factor. This is desirable, since the shadows of supermassive black holes are directly observable [4], and large deviations from general relativity are likely ruled out at that scale.
Last, one may wonder if these deviations, if detected, constitute modified, classical gravity, or contain traces of quantum aspects of gravity. The appearance of $GM$ in the regulator $L$, even in the case of the unmodified Hayward metric, is necessitated by the limiting curvature condition. This condition states that irrespective of its mass, the maximally allowed curvature of an object cannot exceed the Planckian curvature $\mathcal{R}\sim 1/\ell_{\text{Planck}}^{2}$. In other words, the maximum curvature needs to be independent of the object’s mass. This condition (sometimes also called conjecture) is rooted in quantum-gravitational considerations. Therefore, regulators with an unconventional mass dependence, like the ones described in this Letter, may very well one day traced back to a theory of quantum gravity.
Acknowledgements.—We are grateful for support by the National Science Foundation under grant no. PHY-2112460.